# LV-Eval: A Balanced Long-Context Benchmark with

### v1

@@ 3.3 Metric Design
Specifically, to evaluate an answer $A^{\prime}$, we first calculate the recall of several “answer keywords” in $A^{\prime}$. When the recall exceeds a certain threshold (0.2 for Chinese dataset, 0.4 for English datasets), we calculate the F1 score between $A^{\prime}$ and GT answer $A$ as the final score for $A^{\prime}$. otherwise, $A^{\prime}$ gets a zero score. We manually annotate the answer keywords in the GT answer $A$ for hotpotwikiqa-mixup, lic-mixup, loogle-CR-mixup, loogle-MR-mixup, loogle-SD-mixup, multifieldqa-en-mixup, and multifieldqa-zh-mixup. Figure 4 (a) shows an example, demonstrating how this two-stage calculation helps avoid some inflated high evaluation scores.
When calculating the F1 score between $A^{\prime}$ and $A$ in the second stage, we exclude common but non-informative words like ‘the,’ ‘a’, ‘of’, and so on. The word blacklist is constructed as follows. We first summarized the word counts in the generations of Llama2-7B-Chat-hf and ChatGLM3-6B-32K on all datasets and chose the top 100 words that matched the GT answer most frequently. Then, we manually annotate the non-informative words from the 100 words to construct the blacklist. Figure 4 (Lower) shows an example of how the word blacklist aids in calibrating the evaluation scores.
@@ 4 Evaluation
4.1 Evaluation Setup
Models.
We evaluate 2 commercial and 8 open-source LLMs on LV-Eval. Their information is summarized in Table 3.
Figure 5: Average scores across all datasets of 10 LLMs at 5 length levels.
Figure 6: Average scores across all length levels of 10 LLMs on 5 types of tasks.
Figure 7: Average scores across all length levels of 10 LLMs on single-hop QA datasets.
Figure 8: Average scores across all length levels of 10 LLMs on multi-hop QA datasets.
Inference.
We follow the official implementation of all LLMs to conduct their inferences. Greedy sampling is used for generating tokens. For LLMs with a context window size smaller than the length of the data context, we truncate the data context in the middle, and concatenate the head and the tail of the context as input, ensuring that the QA instructions are fully contained within the input.
@@ 4.2 Compare LLMs on LV-Eval
Figure 5 shows the average scores across all 11 datasets of 10 LLMs at different length levels. We can see that (i) Commercial models do not always perform better than open-source models. For instance, ChatGLM3-6B-32k attains the highest accuracy on all length levels. (ii) Models exhibit distinct score trends. From the average scores in Figure 5 and the task-specific scores in Table A2, we can see that the model with the largest context window size, Yi-6B-200k, exhibits the slowest decline of performance from $16k$ to $128k$. For example, its scores at the length level $16k$ is lower than GPT-3.5-16k and GPT-4-8k. Nevertheless, as the length of input context increases, Yi-6B-200k retains a higher score than these two commercial models that need to truncate the input context.
@@
Figure 6 shows the average scores across all 5 length levels of 10 LLMs on 5 types of tasks. We can see that (i) LLMs attain lower scores on multi-hop QA tasks compared to single-hop QA tasks. (ii) Confusing facts insertion adds complexity to the tasks, particularly evident in single-hop QA and single-hop confusion QA.
The detailed results on each dataset of the single-hop QA task type and multi-hop QA task type are shown in Figure 7 and Figure 8, respectively. We can see that (i) Among the multi-hop QA datasets, loogle-CR-mixup and loogle-MR-mixup are particularly challenging. Future research is needed to improve the ability to aggregate multiple pieces of evidence from a long context with distracting and confusing facts. (ii) For single-hop QA datasets, as expected, LLMs can achieve higher scores on datasets without confusing fact insertion, including loogle-SD-mixup and cmrc-mixup. (iii) Several LLMs, namely ChatGLM3-6B-32k, BlueLM-7B-32k-Chat, Yi-6B-200k, and Llama2-7B-32k-Instruct, can achieve relatively high scores on factrecall-en. This indicates that the “needle in a haystack” task might not be challenging enough, emphasizing the need to evaluate LLMs on other tasks, particularly multi-hop QA datasets. (iv) The performance gap between LLMs on factrecall-en and factrecall-zh is especially large, and some open-source LLMs with relatively small context sizes, namely Llama2-7B-Chat-hf (4k context window size), Qwen-7B-8k-Chat, and Vicuna-7B-16k-v1.5, even get near-zero scores. (v) A few LLMs have unbalanced performances on Chinese and English datasets, as illustrated by the results on multifieldqa-en-mixup and multifieldqa-zh-mixup.
Figure 9: Ablation results of the “needle in a haystack” task on ChatGLM3-6B-32k. (a) The results on factrecall-en. (b) The results on factrecall-zh. In each of (a)(b), the top-left, top-right, bottom-left, and bottom-right sub-figures show the results of w.o. “CFI and KPR”, “w. CFI only”, “w. KPR only”, and “w. both CFI and KPR”, respectively. These results illustrate that CFI and KPR are effective in improving the task difficulty.
4.3 Ablation Study of LV-Eval
@@
LongChat-7B-32k-v1.5 direct (w. KPR) 3.84
direct (w.o. KPR) 8.20
w. context (w. KPR) 11.57 10.71 4.77 5.49 2.37
Llama2-7B-32k-Instruct direct (w. KPR) 0.76
direct (w.o. KPR) 2.48
w. context (w. KPR) 3.54 2.31 2.20 1.86 1.62
@@
Table 4: Ablation results for keyword and phrase replacement (KPR): Comparison of “direct (w. KPR)” (applying KPR and direct querying without context), “direct (w.o. KPR)” (direct querying without context), “w. context (w. KPR)” (applying KPR and querying with context). Note that there is only one score result in the first two rows in each section of the table, since the test results of direct querying without context do NOT depend on the context length.
Confusing facts insertion.
Table 6, 7, and 8 show the scores of multiple LLMs on dataset with and without confusing fact insertion. We can see that (i) On multifieldqa-en-mixup and multifieldqa-zh-mixup, the insertion of confusing facts leads to a notable degradation in the scores of LLMs. However, the introduction of confusing facts in the hotpotwikiqa-mixup dataset does not result in severe degradation. (ii) Table 7 and 8 show that the strongest model, ChatGLM3-6B-32k, exhibits the most substantial score degradation on data with confusing fact insertion. For instance, the score of ChatGLM3-6B-32k degrades from $41.46$ to $31.97$ (a degradation of $9.49$) on the $16k$ length level of multifieldqa-en-mixup, while the score degradation of other 5 LLMs falls within the range $\left[0.47,4.89\right]$. This observation suggests that current powerful LLMs may even be more susceptible to confusing information in the context. Future research is needed to enhance the models’ ability to discern information that appears similar but is in fact unrelated. (iii) As the length of the input context increases, the score degradation becomes smaller. This phenomenon can be attributed to two factors: the truncation of confusing facts and a decrease in baseline performance.
Keyword and phrase replacement.
The technique of keyword and phrase replacement aims to eliminate the knowledge leakage and common-sense memorization of LLMs. Intuitively, for datasets sourced from Wikipedia and other widely used corpus, the risk of knowledge leakage is higher. From the results in Table 6, 7, and 8, we observe that: (i) Keyword and phrase replacement brings notable degradation of LLM scores on these three datasets suggesting that knowledge leakage exists in open-source corpus and can be mitigated by the keyword and phrase replacement. (ii) The extent of degradation is relatively consistent across different length levels.
We conduct an additional experiment to illustrate the knowledge leakage issue and the impact of keyword and phrase replacement in Table 4. Specifically, we compare three settings: (i) Directly querying the LLMs to answer the question without the context (“direct (w.o. KPR)”). (ii) Applying keyword and phrase replacement to the QA pair, and directly querying the LLMs without the context (“direct (w. KPR)”). (iii) Applying keyword and phrase replacement to the QA pair and the context, and querying the LLMs to answer the question with the context (“w. context (w. KPR)”).
Table 4 shows that without keyword and phrase replacement, some LLMs can achieve a considerable score even without context. For instance, Yi-6B-200k and ChatGLM3-6B-32k achieve scores of 16.11 and 12.24, respectively, through memorization or common-sense knowledge. Applying keyword and phrase replacement effectively reduces the score without context (6.06 for Yi-6B-200k and 4.96 for ChatGLM3-6B-32k). This helps mitigate the influence of memorization or common-sense knowledge on the assessment of long-context understanding ability.
@@
The factrecall-en and factrecall-zh datasets are constructed to evaluate the “needle in a haystack” [39] ability. We show the ablation results of confusing fact insertion (CFI) and keyword and phrase replacement (KPR) in Figure 9 and Table 9. From the top-left sub-figure in Figure 9, we can see that ChatGLM3-6B-32k attains high accuracy on datasets without CFI and KPR, as long as the input context length is within its context size (32k). However, when either CFI (top-right sub-figure) or KPR (bottom-left sub-figure) is applied, the retrieval accuracy decreases. The accuracy experiences a more severe degradation when both CFI and KPR are applied, particularly evident in factrecall-zh, where a performance collapse is observed. This indicates that there is room for improvement in the model’s ability to accurately identify a specific piece of information from a long context in the presence of interference.
@@
For a given length level $L_{d}$ of the dataset, if the single key information is uniformly distributed in the context, an LLM with a context window size $L_{m}$ can only observe the key information for approximately $\frac{L_{m}}{L_{d}}$ of the time. Thanks to our keyword and phrase replacement technique, we can expect that the LLM cannot get the correct answer through memorization or common sense. Then, ideally, we would not expect to see a metric score much higher than $\frac{L_{m}}{L_{d}}$. However, as shown in Table 5, when using the original F1 metric, due to the undesired matching of non-keywords and non-informative words, the metric score can be a lot higher than $\frac{L_{m}}{L_{d}}$. For instance, ChatGLM3-6B-32k achieves a score of $26.43\%$ on the $256k$ length level of the cmrc-mixup dataset, which significantly exceeds $\frac{L_{m}}{L_{d}}=12.5\%$. Fortunately, our keyword-recall-based metric with the word blacklist returns a score that aligns more closely with human expectations and is more reasonable.
5 Conclusion
In this paper, we introduce LV-Eval, a long-context benchmark with 5 word length levels up to $256k$. The long context size and balanced design of multiple length levels provide a more controllable and comprehensive evaluation of the long-context capability of LLMs. In addition, the construction of LV-Eval involves (i) the incorporation of distraction and confusion to make the benchmark more challenging, (ii) keyword and phrase replacement to mitigate knowledge leakage, and (iii) keyword-recall-based metric for more objective scoring. Ablation studies are conducted to demonstrate the effectiveness of these techniques in benchmark construction.
Furthermore, we assess the performance of 10 state-of-the-art LLMs on LV-Eval, providing insights into their strengths and limitations. We anticipate that LV-Eval will serve as a valuable resource for supporting future research on long-context LLMs.
Acknowledgments
We thank for the insightful discussion with Kai Chen and Songyang Zhang from Shanghai Artificial Intelligence Laboratory. Thanks to all of the annotators who contribute to the project.
@@ Acknowledgments
reference $\frac{L_{m}}{L_{d}}$ (theoretical max score) 100 100 50.00 25.00 12.50

### v2

@@ 3.3 Metric Design
Specifically, to evaluate an answer $A^{\prime}$, we first calculate the recall of several “answer keywords” in $A^{\prime}$. When the recall exceeds a certain threshold (0.2 for Chinese dataset, 0.4 for English datasets), we calculate the F1 score between $A^{\prime}$ and GT answer $A$ as the final score for $A^{\prime}$. otherwise, $A^{\prime}$ gets a zero score. We manually annotate the answer keywords in the GT answer $A$ for hotpotwikiqa-mixup, lic-mixup, loogle-CR-mixup, loogle-MR-mixup, loogle-SD-mixup, multifieldqa-en-mixup, and multifieldqa-zh-mixup. Figure A6 (a) shows an example, demonstrating how this two-stage calculation helps avoid some inflated high evaluation scores.
When calculating the F1 score between $A^{\prime}$ and $A$ in the second stage, we exclude common but non-informative words like ‘the,’ ‘a’, ‘of’, and so on. The word blacklist is constructed as follows. We first summarized the word counts in the generations of Llama2-7B-Chat-hf and ChatGLM3-6B-32K on all datasets and chose the top 100 words that matched the GT answer most frequently. Then, we manually annotate the non-informative words from the 100 words to construct the blacklist. Figure A6 (b) shows an example of how the word blacklist aids in calibrating the evaluation scores.
@@ 4 Evaluation
Models and Inference.
We evaluate 3 commercial and 12 open-source LLMs on LV-Eval. Their information is summarized in Table 3. We follow the official implementation of all LLMs to conduct their inferences. Greedy sampling is used for generating tokens. For LLMs with a context window size smaller than the length of the data context, we truncate the data context in the middle, and concatenate the head and the tail of the context as input, ensuring that the QA instructions are fully contained within the input.
Model Name SFT Context Length HuggingFace / API Endpoint
Llama2-7B-Chat-hf Touvron et al. (2023) $4k$ meta-llama/Llama-2-7b-chat-hf
Qwen-7B-8k-Chat Bai et al. (2023a) $8k$ Qwen/Qwen-7B-Chat
Llama3-8B-Instruct AI@Meta (2024) $8k$ meta-llama/Meta-Llama-3-8B-Instruct
Vicuna-7B-16k-v1.5 Zheng et al. (2023) $16k$ lmsys/vicuna-7b-v1.5-16k
ChatGLM3-6B-32k Zeng et al. (2022) $32k$ THUDM/chatglm3-6b-32k
BlueLM-7B-32k-Chat Team (2023) $32k$ vivo-ai/BlueLM-7B-Chat-32K
LongChat-7B-32k-v1.5 togetherAI (2023) $32k$ lmsys/longchat-7b-v1.5-32k
Qwen2.5-72B-Instruct-128k Yang et al. (2024) $128k$ Qwen/Qwen2.5-72B-Instruct
Meta-Llama-3.1-70B-128k-Instruct Meta (2024) $128k$ meta-llama/Llama-3.1-70B-Instruct
Yi-6B-200k Yi (2023) $200k$ 01-ai/Yi-6B-200K
Llama3-8B-1M gradient.ai (2024) $1048k$ gradientai/Llama-3-8B-Instruct-Gradient-1048k
GPT-4-8k Achiam et al. (2023) $8k$ gpt-4-0613
GPT-3.5-16k Ye et al. (2023) $16k$ gpt-3.5-turbo-1106
Moonshot-V1-128k moonshot (2024) $128k$ moonshot-V1-128k
Table 3: Information of evaluated LLMs.
@@ 4.1 Compare LLMs on LV-Eval
Figure 4: Overall results on different length levels and types of datasets. (a) Average scores across all datasets of 15 LLMs at 5 length levels. The circle markers represent open-source models of 6-9B; The square markers represent larger open-source models around 70B; The triangle markers represent commercial APIs. Note that we only evaluate Moonshot-v1-128k on $16k$ and $32k$ length levels due to the high cost. (b) Average scores across all length levels of 14 LLMs on 5 types of datasets. “CQA” refers to QA datasets with CFI.
@@
Figure 4 (a) shows the average scores across all 11 datasets of 15 LLMs at different length levels. We can see that (i) The commercial Moonshot-v1 and recent large-scale open-source models (Qwen2.5-72B, Llama-3.1-70B) achieve the best performances. Notably, both Qwen2.5-72B and Moonshot-v1 obtain average scores exceeding 40 at $16k$ and $32k$ lengths. (ii) Among open-source models with parameter sizes in the 6-9B range, GLM-4-9B achieves the best performance, even outperforming Llama-3.1-70B on longer lengths ($128k$, $256k$). (iii) Models exhibit distinct score trends, resulting in different relative rankings across different length levels. For example, the model with the largest context window size, Llama3-8B-1M, exhibits one of the slowest decline of performance from $16k$ to $128k$. Specifically, its scores at the length level $16k$ is lower than ChatGLM3-6B-32k and BlueLM-7B-32k-Chat. Nevertheless, as the length of input context increases, Llama3-8B-1M retains a higher score than these two models that need to truncate the input context. The similar phenomenon can be observed between Yi-6B-200k and two GPTs, as well as between GLM-4-9B and Llama-3.1-70B. (iv) Models exhibit a sharp performance drop, often occurring prior to or when the context length exceeds their supported context length. Some models experience a sharp decline in performance once the context length exceeds their supported context length. For instance, Qwen2.5-72B-128k, GLM-4-9B-128k, and ChatGLM3-6B-32k exhibit a sharp performance drop after $128k$, $128k$, and $32k$, respectively. In contrast, some other models encounter this sharp performance drop earlier. For instance, Llama-3.1-70B-128k exhibit a sharp performance drop after $64k$, despite its stated support for $128k$ context.
@@
Figure 4 (b) shows the average scores across all 5 length levels on 5 task types. We can see that (i) LLMs attain lower scores on multi-hop QA tasks compared to single-hop QA tasks. (ii) Confusing facts insertion adds complexity to the tasks, particularly evident in single-hop QA and single-hop confusion QA. See Appendix C for more detailed results and Appendix F for example failure cases.
4.2 Ablation Study of LV-Eval Techniques
Confusing facts insertion.
Table A8, A9, and A10 show the scores of multiple LLMs on dataset with and without CFI. We can see that (i) On multifieldqa-en-mixup and multifieldqa-zh-mixup, CFI leads to a notable degradation in the scores of LLMs. However, CFI in the hotpotwikiqa-mixup dataset does not result in severe degradation. (ii) Table A9 and A10 show that a strong model, ChatGLM3-6B-32k, exhibits the most substantial score degradation on data with CFI. For instance, the score of ChatGLM3-6B-32k degrades from $41.46$ to $31.97$ (a degradation of $9.49$) on the $16k$ length level of multifieldqa-en-mixup, while the score degradation of other 5 LLMs falls within the range $\left[0.47,4.89\right]$. This observation suggests that current powerful LLMs may even be more susceptible to confusing information in the context. Future research is needed to enhance the models’ ability to discern information that appears similar but is in fact unrelated. (iii) As the length of the input context increases, the score degradation becomes smaller. This phenomenon can be attributed to two factors: the truncation of confusing facts and a decrease in baseline performance.
Keyword and phrase replacement.
The technique of KPR aims to eliminate the knowledge leakage and common-sense memorization of LLMs. Intuitively, for datasets sourced from Wikipedia and other widely used corpus, the risk of knowledge leakage is higher. From the results in Table A8, A9, and A10, we observe that: (i) KPR brings notable degradation of LLM scores on these three datasets suggesting that knowledge leakage exists in open-source corpus and can be mitigated by KPR. (ii) The extent of degradation is relatively consistent across different length levels.
We illustrate the knowledge leakage issue and the impact of KPR in Table 4. Specifically, we compare three settings: (i) Directly querying the LLMs to answer the question without the context (“direct (w.o. KPR)”). (ii) Applying KPR to the QA pair, and directly querying the LLMs without the context (“direct (w. KPR)”). (iii) Applying KPR to the QA pair and the context, and querying the LLMs to answer the question with the context (“w. context (w. KPR)”).
Table 4 shows that without KPR, some LLMs can achieve a considerable score even without context. For instance, Yi-6B-200k and ChatGLM3-6B-32k achieve scores of 16.11 and 12.24, respectively, through memorization or common-sense knowledge. Applying KPR decreases the score without context (6.06 for Yi-6B-200k and 4.96 for ChatGLM3-6B-32k). This helps mitigate the influence of memorization or common-sense knowledge on the assessment of long-context understanding ability.
@@ 4.2 Ablation Study of LV-Eval Techniques
Table 4: Ablation results for KPR. “direct (w. KPR)”: Apply KPR and direct query without context; “direct (w.o. KPR)”: Direct query without context; “w. context (w. KPR)”: Apply KPR and query with context. Note that there is only one result in the first two rows in each section of the table, since the results of direct querying without context do NOT depend on the context length.
@@
The factrecall-en and factrecall-zh datasets are constructed to evaluate the enhanced NIAH (Kamradt, 2023) ability. The traditional NIAH evaluation is basically a retrieval task, asking LLMs to find the answer or passkey in long context, which is too simple for majority of LLMs that they can easily get high scores after task oriented training. Therefore we enhance the NIAH evaluation with CFI and KPR to assess LLM’s positional consistency of retrieval while challenging their comprehension and anti-interference abilility.
We show the ablation results of CFI and KPR in Figure 5 and Table A11. From the first column in Figure 5, we can see that ChatGLM3-6B-32k attains high accuracy on datasets without CFI and KPR, as long as the input context length is within its context size (32k). However, when either CFI (second column of sub-figure) or KPR (third column sub-figure) is applied, the retrieval accuracy decreases. The accuracy experiences a more severe degradation when both CFI and KPR are applied, particularly evident in factrecall-zh, where a performance collapse is observed (See Appendix F for an example failure case). This indicates that there is room for improvement in the model’s ability to accurately identify a specific piece of information from a long context in the presence of interference.
Figure 5: Ablation results of the “needle in a haystack” task on ChatGLM3-6B-32k. (a) factrecall-en. (b) factrecall-zh. In each of (a)(b), from left to right, the four sub-figures show the results of w.o. “CFI and KPR”, “w. CFI only”, “w. KPR only”, and “w. both CFI and KPR”, respectively. These results illustrate that CFI and KPR are effective in improving the task difficulty.
@@
For a given length level $L_{d}$ of the dataset, if the single key information is uniformly distributed in the context, an LLM with a context window size $L_{m}$ can only observe the key information for approximately $\frac{L_{m}}{L_{d}}$ of the time. Thanks to our KPR technique, we can reasonably expect that the LLM cannot get the correct answer through memorization or common-sense knowledge. Furthermore, in our free-form QA tasks, unlike in multiple-choice settings, the LLM cannot easily guess the correct answer. Therefore, we would not expect to see a metric score much higher than $\frac{L_{m}}{L_{d}}$. In other words, $\frac{L_{m}}{L_{d}}$ is the ideal score assuming that all answers are correctly extracted from and only from the context, and the metric accurately measures the answer correctness.
However, as shown in Table 5, when using the original F1 metric, the metric score can be significantly higher than $\frac{L_{m}}{L_{d}}$ due to the undesired matching of non-keywords and non-informative words. For example, ChatGLM3-6B-32k achieves a score of $26.43\%$ on the $256k$ length level of the cmrc-mixup dataset, which greatly exceeds $\frac{L_{m}}{L_{d}}=12.5\%$. In contrast, our keyword-recall-based metric with the word blacklist is a more meaningful metric, as its scores are more aligned with our expectations (i.e., smaller than $\frac{L_{m}}{L_{d}}$).
@@
reference $\frac{L_{m}}{L_{d}}$ 100 100 50.00 25.00 12.50